# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 12 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.217217217See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@ 12.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.218218218The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.219219219This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.220220220A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:
@@ 12.6 Simulation Study: Confounded Retail Pricing MDP
The action affects the next state only through the mediator: $P(s{+}1\mid s,M_{t},Z_{t})$ depends on $M_{t}$ and $Z_{t}$ but not on $A_{t}$ or $U_{t}$ directly. This enables all four identification strategies simultaneously: $Z_{t}$ satisfies the backdoor criterion, $M_{t}$ satisfies the front-door criterion, $\text{IV}_{t}$ satisfies relevance and exclusion, and $W_{t}^{(1)},W_{t}^{(2)}$ satisfy the proximal conditions.224224224Rewards are r(s,a)=1r(s,a)=-1 for $s<4$, $\gamma=0.9$. The target policy always promotes. The true interventional transition probability is $P(s{+}1\mid s,\operatorname{do}(\text{promote}))=0.615$.
I compare six estimators: oracle, naive (biased per Lemma L1), backdoor (Equation 119), front-door (Equation 120), Wald IV (Equation 122), and proximal (Equations 123124).225225225Each configuration uses 2,000 trajectories averaged over 20 seeds.
@@ 13.2 How Reinforcement Learning Advances Economics
When RL-trained agents are actually deployed in markets, they become economic actors subject to empirical scrutiny. Algorithmic pricing agents, for instance, have been shown to sustain supra-competitive prices through reward-based learning without explicit communication. This is not a speculative concern; competition authorities in the EU, US, and UK are actively investigating whether algorithmic coordination constitutes tacit collusion.226226226The companion thesis (Rawat, 2026) develops a framework for evaluating algorithmic inefficiency and collusion risk in algorithmically mediated markets, combining simulators with factorial experimental designs. As algorithmic agents proliferate in pricing, trading, content recommendation, and resource allocation, the empirical study of machine behavior using economic tools becomes an increasingly important research program.
@@ 13.4 Conclusion
Reinforcement learning is a welcome addition to the economist’s toolkit but is not a replacement for existing methods. It extends the reach of dynamic programming to problems that were previously intractable, and it connects naturally to econometric frameworks through shared mathematical foundations. The research agendas outlined above, from building economic simulators to formalizing the role of structural assumptions in sample complexity to developing inference procedures for learned policies, are concrete and tractable. Progress on them will require sustained collaboration between the two fields, drawing on economics for “structure” and institutional context and on RL for computation.

### 23 Mar 2026 v4

@@ 12 Reinforcement Learning and Causal Inference
Reinforcement learning algorithms solve Markov decision processes by estimating value functions or optimizing policies from sampled transitions. Two assumptions underlie the standard formulation. First, the agent’s action at time $t$ is determined solely by the observed state $s_{t}$ and the agent’s policy $\pi(a\mid s_{t})$; there is no unobserved variable (confounder) simultaneously influencing both the action and the reward or state transition.227227227See Section 2 for the terminological mapping between “outcome” in causal inference and the corresponding RL quantities. Throughout this chapter, I reserve “outcome” for its causal inference meaning and use the specific RL quantity (reward, state, return, value) elsewhere. Second, the observed state $s_{t}$ is sufficient for prediction, so that conditioning on $s_{t}$ renders future states independent of past history. Both assumptions are the Markov property restated in causal language. Both fail routinely in economic applications.
@@ 12.2 The Confounded MDP
When unobserved confounders influence both the behavioral policy and the transitions or rewards, the MDP is confounded. This formalization, developed by Zhang and Bareinboim (2019), Zhang and Bareinboim (2020), and Kallus and Zhou (2020), provides the foundation for causal reasoning in sequential decision problems. The key tool is Pearl’s do-operator. The interventional distribution $P(Y\mid\operatorname{do}(X=x))$ is the distribution that arises when $X$ is set externally rather than observed passively, severing all incoming causal influences on $X$ while leaving the remaining data-generating process intact.228228228The do-operator is formalized within the structural causal model (SCM) framework of Pearl (2009). An SCM specifies endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, structural equations $V_{i}=f_{i}(\text{pa}(V_{i}),U_{i})$, and a distribution $P(\mathbf{U})$. The intervention $\operatorname{do}(X=x)$ replaces the structural equation for $X$ with a constant, producing the interventional distribution. See Pearl (2009) for the complete framework, including the causal hierarchy (association, intervention, counterfactual) and general identification theory.
@@
The importance-sampling estimator $\hat{V}^{\pi}_{\text{IS}}=\frac{1}{N}\sum_{i=1}^{N}\prod_{t=0}^{T}\frac{\pi(a_{t}^{(i)}\mid s_{t}^{(i)})}{\mu(a_{t}^{(i)}\mid s_{t}^{(i)})}G^{(i)}$, where $G^{(i)}=\sum_{t=0}^{T}\gamma^{t}R_{t}^{(i)}$ is the discounted return of trajectory $i$ and $T$ is the trajectory length, is also biased because the propensity $\mu(a\mid s)$ is not the true behavioral propensity $\mu(a\mid s,u)$.229229229This is the sequential analogue of the omitted variable bias in linear regression. In the static case, regressing $Y$ on $X$ without controlling for a confounder $U$ yields a biased coefficient. In the sequential case, the bias propagates through the Bellman recursion and can amplify over the horizon.
@@
Suppose a set of observed variables $\mathbf{Z}_{t}$ satisfies the backdoor criterion relative to $(A_{t},S_{t+1})$ in the causal graph of the confounded MDP, meaning that $\mathbf{Z}_{t}$ blocks all backdoor paths from $A_{t}$ to $S_{t+1}$ and no element of $\mathbf{Z}_{t}$ is a descendant of $A_{t}$.230230230A backdoor path from $A_{t}$ to $S_{t+1}$ is any path in the causal graph that begins with an arrow into $A_{t}$ (i.e., a non-causal path). In the confounded MDP, $A_{t}\leftarrow U_{t}\rightarrow S_{t+1}$ is a backdoor path: $U_{t}$ causes both $A_{t}$ and $S_{t+1}$, creating a spurious association. Blocking all such paths by conditioning on appropriate variables eliminates the confounding bias. See Pearl (2009), Chapter 3. Then the interventional transition probability is identified:
@@ 12.6 Simulation Study: Confounded Retail Pricing MDP
The action affects the next state only through the mediator: $P(s{+}1\mid s,M_{t},Z_{t})$ depends on $M_{t}$ and $Z_{t}$ but not on $A_{t}$ or $U_{t}$ directly. This enables all four identification strategies simultaneously: $Z_{t}$ satisfies the backdoor criterion, $M_{t}$ satisfies the front-door criterion, $\text{IV}_{t}$ satisfies relevance and exclusion, and $W_{t}^{(1)},W_{t}^{(2)}$ satisfy the proximal conditions.234234234Rewards are r(s,a)=1r(s,a)=-1 for $s<4$, $\gamma=0.9$. The target policy always promotes. The true interventional transition probability is $P(s{+}1\mid s,\operatorname{do}(\text{promote}))=0.615$.
I compare six estimators: oracle, naive (biased per Lemma L1), backdoor (Equation 124), front-door (Equation 125), Wald IV (Equation 127), and proximal (Equations 128129).235235235Each configuration uses 2,000 trajectories averaged over 20 seeds.
@@ 13.2 How Reinforcement Learning Advances Economics
When RL-trained agents are actually deployed in markets, they become economic actors subject to empirical scrutiny. Algorithmic pricing agents, for instance, have been shown to sustain supra-competitive prices through reward-based learning without explicit communication. This is not a speculative concern; competition authorities in the EU, US, and UK are actively investigating whether algorithmic coordination constitutes tacit collusion.236236236The companion thesis (Rawat, 2026) develops a framework for evaluating algorithmic inefficiency and collusion risk in algorithmically mediated markets, combining simulators with factorial experimental designs. As algorithmic agents proliferate in pricing, trading, content recommendation, and resource allocation, the empirical study of machine behavior using economic tools becomes an increasingly important research program.
@@ 13.4 Conclusion
Reinforcement learning is a welcome addition to the researcher’s toolkit but is not a replacement for existing methods. It extends the reach of dynamic programming to problems that were previously intractable, and it connects naturally to econometric frameworks through shared mathematical foundations. The research agendas outlined above, from building economic simulators to formalizing the role of structural assumptions in sample complexity to developing inference procedures for learned policies, are concrete and tractable. Progress on them will require sustained collaboration between the two fields, drawing on economics for “structure” and institutional context and on RL for computation.